# Comparison of two Cooperative Maneuver Planning Approaches at a Real-World T-Junction

### v2

@@ 7 REAL-WORLD EXPERIMENTS
To support our simulative results, we also conducted test drives in real-world traffic. Although a similar quantitative evaluation as in simulation is not possible, this shows that our cooperative maneuver planning system can be successfully deployed to a real-world setup of connected intelligent infrastructure and CAVs.
7.1 Pilot Site Ulm-Lehr
@@
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [1]. An aerial view of the site is shown in Fig. 9. The unsignalized intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. An edge server resides in the same 5G network cell and a fusion and tracking module [42] combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules. The latter also run on the edge server. We employed three CAVs, of which one used the trajectory planning from [5] and the other two implemented the concept from [43].
@@
The scenario that was investigated in real-world experiments comprises the three CAVs approaching the intersection from each of its accesses, as depicted in Fig. 1. Vehicle approaches the intersection on the main road coming from north and following the road. According to ordinary priority rules, this vehicle has the right of way. Vehicle starts at the other main road access (east) and intends to turn left at the intersection. This vehicle would have to yield to the oncoming traffic but not to any vehicle coming from the side road. With vehicle approaching the intersection from the west on the minor road, it would have to yield to all cross traffic. As it intends to turn right at the intersection, there is no conflict with but only with . Figure 1 illustrates a typical cooperative maneuver. If vehicle waives its right of way and possibly slows down while approaching, it allows both other CAVs to cross the intersection virtually simultaneously. Depending on the timing, vehicles and can simply traverse the intersection unimpeded. Ideally, vehicle should not be required to stop, but slow down smoothly and accelerate again once the intersection is cleared.
7.2 Evaluation in Real-World
@@
To obtain a reliable quantitative evaluation in real traffic, the initial scenario depicted in Fig. 1 was driven numerous times. Thereof, five baseline runs have been performed by driving according to priority rules, twelve runs under the command of the optimization-based planner, and 20 runs employing the RL planner. The cooperative maneuvers have been planned reactively with respect to surrounding traffic. One of the evaluation runs using the RL planner was recorded on video for demonstration purposes and is available at [41]. The vehicles’ trajectories have been recorded and clipped to the common evaluation segment, which was also used for the simulative evaluation. Because all experiments have been conducted in real traffic at a public intersection, the initial conditions are not ensured to match perfectly. While the safety drivers guided the test vehicles carefully to the engagement point, minor deviations in timing and initial speed remain. During all maneuver runs that are contained in the evaluation, the automated driving system was successfully engaged for at least the whole evaluation segment.
@@
8 CONCLUSION
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been evaluated. The proposed approaches have been compared in a realistic simulation and in real-world experiments using three prototype CAVs. Compared to HDVs adhering to priority rules and non-cooperative CAVs, the cooperative maneuvers lead to improved traffic efficiency. In simulated fully automated traffic, a significant reduction of delay due to interaction and of the share of stopped vehicles has been shown. Even in mixed traffic with only CAV penetration, the proposed planners still found a viable cooperative maneuver in about one third of the scenarios, despite not relying on any cooperation or turn direction information from the HDVs. Both cooperative planners demonstrated their real-world applicability in public traffic and yielded a significant reduction in delay due to interaction. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven connected vehicles and vulnerable road users.

### v5

@@ VI Real-World Experiments
To support our simulative results, we also conducted test drives in real-world traffic. Although a similar comparative evaluation as in simulation is not possible due to fluctuating real-world traffic conditions, this shows that our cooperative maneuver planning system can be successfully deployed to a real-world setup of connected intelligent infrastructure and CAVs.
VI-A Pilot Site Ulm-Lehr
@@
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [4]. An aerial view of the site is shown in Fig. 7. The unsignalized intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. A fusion and tracking module [13] on an edge server combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules on the same edge server. We employed three CAVs, one of which used the trajectory planning from [32] while the other two implemented the concept from [38]. The CAVs, infrastructure, and edge server are connected via a 5G SA cellular network with a communication latency of about 15 ms.
@@
The scenario that was investigated in real-world experiments comprises the three CAVs approaching the intersection from each of its accesses, as depicted in Fig. 1. Vehicle approaches the intersection on the main road coming from north and following the road. According to ordinary priority rules, this vehicle has the right of way. Vehicle starts at the other main road access (east) and intends to turn left at the intersection. This vehicle would have to yield to the oncoming traffic but not to any vehicle coming from the subordinate road. With vehicle approaching the intersection from the west on the subordinate road, it would have to yield to all cross traffic. As it intends to turn right at the intersection, there is no conflict with but only with . Fig. 1 illustrates a typical cooperative maneuver. If vehicle waives its right of way and possibly slows down while approaching, it allows both other CAVs to cross the intersection virtually simultaneously. Depending on the timing, vehicles and can simply traverse the intersection unimpeded. Ideally, vehicle should not be required to stop, but slow down smoothly and accelerate again once the intersection is cleared.
VI-B Evaluation in Real-World
@@
To obtain a reliable quantitative evaluation in real traffic, the initial scenario depicted in Fig. 1 was driven numerous times. Thereof, five baseline runs have been performed by driving according to priority rules, twelve runs under the command of the optimization-based planner, and 20 runs employing the RL planner. The cooperative maneuvers have been planned reactively with respect to surrounding traffic, while the CAVs approach the intersection. This manifests in alternative maneuvers being proposed if the original proposal turns out to be less effective or even infeasible. One of the evaluation runs using the RL planner was recorded on video for demonstration purposes and is available at [31]. The vehicle trajectories have been recorded and clipped to the common evaluation segment, which was also used for the simulative evaluation. Because all experiments have been conducted in real traffic at a public intersection, each run had slightly different initial conditions. While the safety drivers guided the test vehicles carefully to the engagement point, minor deviations in timing and initial speed remain. During all maneuver runs that are contained in the evaluation, the automated driving system was successfully engaged.
@@
VII Conclusion
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been evaluated in a realistic simulation and in real-world experiments. Especially in simulated fully automated traffic, cooperative maneuvers achieved a significant improvement of traffic efficiency. Even in simulated mixed traffic with only 50 % CAV penetration, the planning approaches successfully performed cooperative maneuvers in about one third of the scenarios and increased the efficiency, despite not relying on any cooperation from HDVs. Overall, the optimization-based planner yielded higher efficiency gains, while also causing a larger increase of criticality measures. Experiments with three prototype CAVs in public traffic demonstrated the real-world applicability of both cooperative planners and showed a significant efficiency increase. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven connected vehicles and vulnerable road users.